A manufacturer of light bulbs finds that one light bulb model has a
mean life span of 1025 h with a standard deviation of 87 h. What percent of these
light bulbs will last?
a. at least 950 h?
b.
between 800 and 900 h?
1
Expert's answer
2020-12-15T02:21:09-0500

Let "X" be a life span of a bulb: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"

Given "\mu=1025\ h,\sigma=87\ h"

a)

"P(X\geq950)=1-P(X<950)"

"=1-P(Z<\dfrac{950-1025}{87})\approx1-P(Z<0.862069)"

"\approx1-0.194325=0.805675"

b)

"P(800<X<900)=P(X<900)-P(X\leq800)"

"=P(Z<\dfrac{900-1025}{87})-P(Z<\dfrac{800-1025}{87})"

"\approx P(Z<-1.436782)-P(Z<-2.586207)"

"\approx0.075390-0.004852=0.070538"